Fermi National Laboratory


Accelerator Update for March 14, 2005 - March 16, 2005

Monday March 14
The day shift began with Operations monitoring the Antiproton Source (Pbar) stack and the Recycler stash, with the Tevatron (TeV) in shot setup, with MiniBooNE taking beam, and with NuMI and Switchyard (SY) beam off due to shot setup.

Operations established store 4043 at 9:51 AM with an initial luminosity of 94.92E30. Pbar resumed stacking.

Operations established beam to the Meson MCenter experiment at 10:44 AM.

Operations established beam to NuMI at 11:01 AM.

Operations began stashing the Recycler with antiprotons at 3:03 PM. This stash was accomplished with a new transfer scheme that took only forty minutes for four transfers.

At 3:57 PM, a Linac RF station (LRF2) tripped off for the first of four times in fifteen minutes. It reset okay each time.

At 4:45 PM, a Booster RF station (BRF5) tripped off three times in a minute. It reset okay.

At 5:49 PM, LRF2 tripped off for the first of two times in twenty-four minutes. It reset okay.

At 7:37 PM, BRF4 tripped off. It reset okay.

At 8:51 PM, BRF4 began tripping off repeatedly. Operators bypassed the station for about an hour and then turned it back on.

At 9:50 PM, the Linac Buncher tripped off. It reset, eventually.

Tuesday March 15
The midnight shift began with Operations monitoring store 4043 and the stash, with Pbar stacking, and with MiniBooNE, MCenter, and NuMI taking beam.

At 12:34 AM, the Linac Buncher tripped again. It reset okay.

At 5:20 AM, the Linac Buncher tripped again. It reset okay.

At 9:32 AM, BRF4 tripped off. It reset okay.

Pbar and Main Injector (MI) experts began a slipstacking study at 9:46 AM. The study lasted for two hours.

At 10:21 AM, BRF4 tripped off twice. Experts reported that these trips were due to bunch rotation and at present there was no solution to the trips.

TeV experts began an end-of-store study at 12:02 PM. The study ended and then operators terminated store 4043 at 12:42 PM.

Operations put the TeV into shot setup at 12:57 PM.

Preaccelerator experts accessed the H- Source at 2:40 PM to check the generator brushes and accessed the I- Source to replace the source.

Operations established store 4044 at 3:18 PM with an initial luminosity of 100.7E30. Pbar resumed stacking.

NuMI and MCenter resumed taking beam at 3:36 PM.

At 5:11 PM, BRF4 tripped off several times in about twenty minutes. The last time it tripped, the station was bypassed.

At 5:40 PM, the Linac Buncher tripped off. Operators had it reset twelve minutes later.

At 6 PM, BRF5 tripped off repeatedly. Operators contacted an expert who had it back up by 6:08 PM.

At 6:54 PM, a Booster vacuum pump tripped off due to a vacuum burst, which then caused two RF stations to trip off. Everything reset okay.

Operations put the Recycler into shot setup to stash antiprotons at 9:57 PM. Operators completed the transfers by 11:03 PM. During the stash, Operations returned BRF4 to service.

NuMI and MCenter resumed taking beam by11:20 PM.

Wednesday March 16
The midnight shift began with Operations monitoring store 4044 and the stash, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

At 3:34 AM, BRF4 tripped off for the first of three times in two minutes. Operators bypassed the station after the last trip.

At 4:17 AM, the cooling system for the Meson M01 target train tripped off. This disabled the MCenter beam. Operations contacted an expert to come in and make repairs.

Operators turned on BRF4 at 6:12 AM.

The Plan for Wednesday
The plan for today is to leave the present store in until 9 PM. During this time, the run coordinator will allow Pbar to conduct a reverse proton study. After Operations terminates the store at 9 PM, the Collider Detector at Fermilab and D-Zero will be allowed to make an access. Shot setup will begin around midnight.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 3/16/2005   email Fermilab